a particle is projected vertically upward from A with a speed of 50 m/s and another is dropped simultaneously from B which is 200 m vertically above A. They cross each other ?
 

Grade:11

3 Answers

Vishal Kumar
5 years ago
u1= 50m/s & u2= 0
relative speed is u1 - u2 = 50m/s . 
relative separation is 200m. 
time, t= 200/50 = 4s
Lok Nitesh
5 years ago
h/
Lok Nitesh
5 years ago
h/u = 200/50 = 4 sec as if a body thrown 
Vertically upwards from a height 'h’ and if another particle thrown from ground both will meet at t= h/u